OpenStudy (anonymous):

Write an equation of the circle centered at (7, -2) that passes through (-10, 0)

OpenStudy (anonymous):

oh circles :3

OpenStudy (zale101):

to know the radius between the center and the passing point, means that u use the The Distance Formula

OpenStudy (anonymous):

well... you guys are no fun :D I was trying to see if joseph was still among us living folk ^.^ haha

OpenStudy (zale101):

\[d=\sqrt{(x_2-x_1)+(y_2-y_1)^2}\]

OpenStudy (yttrium):

This is very easy. First, solve for its radius. \[r = \sqrt{((-10)-7)^2+(0-(-2)^2)} = \sqrt{(-17)^2+2^2}\]

OpenStudy (anonymous):

also... technically, you don't need r... just r^2... nasty radicals :D \[\large r^2 = (-17)^2 + 2^2\]